What will be the pressure inside the sphere if it is then placed in a mixture of ice and water? Express your answer to two significant figures and include the appropriate units Review A rigid sphere has a valve that can be opened or closed. The sphere with the valve open is placed in boiling water in a room where the air pressure is 1.0 atm. After a long period of time has elapsed the valve is closed P2Value Units Submit Request Answen Part B What will be the pressure inside the sphere if it is then placed in an insulated box filed with dry ice? Express your answer to two significant figures and include the appropriate units
Answer #1

The idea is that pressure is proportional to temperature, if The temp is in Kelvins

    boiling water T =373 K     

p = 1.0 atm

    freezing water    T =273 K  

So,

P2 = 1.0 *273/373

P2=1.0*0.7319

P2=    0.732atm

    dryice T =   195 K  

So,

P2=   1.0 * 195 / 373

P2=1.0*0.5228

P2= 0.523 atm

I hope help you !!